# Reward-Safety Balance in Offline Safe RL via Diffusion Regularization

### v1

@@ 5.3 Ablation Studies
Impact of Different Cost Limits. We evaluate our algorithm’s performance under varying cost limits $l=10,20$, and $30$, analyzing the learned policies’ behavior for each budget setting. The ablation results are presented in Figure 3. Across all cost limit choices, our algorithm consistently achieves zero violations of the safety constraints, demonstrating its robustness to varying cost thresholds. This highlights the adaptability and reliability of our approach in maintaining safety compliance. While COptiDICE also strictly adheres to the cost limits, our method consistently outperforms it in terms of normalized return, demonstrating its superior ability to balance safety and reward optimization. Our main result in Table 1 in evaluated under cost limit $l=10$.
Figure 3: Normalized return and cost under varying cost limits ($l=10,20,30$). Since the cost is normalized relative to the corresponding cost limits, the safety threshold is consistently $1.0$ for all normalized returns across the different cost limits. The dashed line in the figure represents the safety boundary.
@@
Temperature Parameter Setting . We evaluate three temperature ($\beta$) schedules: (a) Constant, where $\beta_{t}=\beta_{0}$ for all timesteps $t$; (b) Square-root growth, where $\beta_{t}=\sqrt{c_{0}+c_{1}t}$, with $c_{0}$ and $c_{1}$ as constants; and (c) Linear growth, where $\beta_{t}=c_{0}+c_{1}t$, where $t$ represents the epoch number, and the same $\beta$ value is maintained within each epoch. Specifically, we compare three beta schemes: $\beta_{t}=0.02$, $\beta_{t}=0.01\sqrt{t}$ and $\beta_{t}=0.01t+0.04$. The comparison between these schemes is illustrated in Figure 4.
@@ 6 Related Work
Diffusion Model for Offline RL. Due to their strong generative capabilities, diffusion models have been extensively employed in offline RL for modeling heterogeneous behavioral datasets. In general, diffusion models have primarily been used for either planning Janner et al. (2022); He et al. (2024); Ajay et al. (2022) or modeling policy distributions Chi et al. (2023); Wang et al. (2022); Fang et al. (2024); Hansen-Estruch et al. (2023); Chen et al. (2023); Lu et al. (2023). However, a significant shortcoming of diffusion models is their slow inference speed, as the reverse denoising process requires many iterations. Various strategies have been proposed to address this issue, such as DDIM Song et al. (2020a), which reduces the sampling steps by sampling on subsequences, and DPM-Solver Lu et al. (2022), which minimizes the sampling steps to around 10 by adopting a specialized diffusion ODE solver.
Safe RL. Existing literature on safe RL predominantly focuses on the online setting (see Gu et al. (2024c) for a comprehensive review). Among these studies, constrained optimization-based methods have garnered significant attention for their strong performance and broad applicability (Achiam et al., 2017; Tessler et al., 2018; Yang et al., 2020b; Liu et al., 2022). For example, Achiam et al. (2017) introduced CPO, which integrates TRPO (Schulman, 2015) into a constrained optimization framework to derive safe policies. Primal-dual methods employing Lagrangian optimization have also gained popularity (Chow et al., 2018; Calian et al., 2020; Ding et al., 2020; Ying et al., 2022; Zhou et al., 2023). These methods dynamically adjust the Lagrange multipliers in response to safety violations, enabling effective management of safety constraints during training. In the category of primal-based approaches, Xu et al. (2021) proposed CRPO, which alleviates the need for tuning the dual variable. Building on CRPO, Gu et al. (2024a) employed gradient manipulation to balance the conflict between reward and cost. However, these online methods often require extensive interactions with the environment and are highly sensitive to critic estimation, making them impractical for real-world applications due to safety concerns, high costs, or limited accessibility during data collection.
Safe Offline RL. Recently, significant attention has been devoted to addressing the safe RL problem in an offline setting, where agents no longer have access to online interactions with the environment. For instance, Xu et al. (2022) proposed learning RL policies from offline data by assigning high cost values to both out-of-distribution (OOD) actions and unsafe actions. Recent works, such as Lin et al. (2023), incorporate the diffuser method (Janner et al., 2022) into the safe setting to learn safe policies. Similarly, Liu et al. (2023b) utilized the decision transformer framework (Chen et al., 2021) to facilitate the learning of safe policies. In another study, Zheng et al. (2024) employed diffusion policies learned through energy-based weighted regression methods (Lu et al., 2023) to ensure adherence to hard constraints.
7 Conclusion
In this work, we address the challenges of offline safe RL, where the primary difficulties arise from the inherent nature of offline datasets and the additional safety constraints. To tackle these challenges, we propose the DRCORL method, which is designed to learn a safe policy with high-reward performance across a wide range of challenging offline safe RL tasks. DRCORL achieves high-speed inference and dynamically resolves conflicts between reward and cost functions, owing to the simple structure of its policy network. We show that our method outperforms existing approaches, which often suffer from either slow inference due to complex policy network architectures or the inability to achieve high returns. Nevertheless, our algorithm can be further studies and improved on various fronts. 1) DRCORL still relies on a pretraining procedure to obtain a diffusion model for imitating the behavioral policy and a critic function that approximates the optimal Q-function. 2) Achieving zero constraint violations remains challenging, particularly under the offline setting. Developing algorithms that are robust to variations in dataset quality remains an important future direction. Despite these limitations, our work provides valuable insights and inspiration for addressing generalization and safety challenges in constrained offline RL.
@@ Appendix A Proofs of Theoretical Results
$\left|\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\mathbb{E}_{a\sim\pi_{b}(\cdot|s)}[A^{\tilde{\pi}}_{c}(s,a)]\right|$ $\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\left|\int_{\mathcal{A}}\pi_{b}(a|s)Q^{\tilde{\pi}}(s,a)-\tilde{\pi}(a|s)Q^{\tilde{\pi}}(s,a)da\right|$ (33)
@@
$\leq\frac{c_{\max}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\int_{\mathcal{A}}|\tilde{\pi}(a|s)-\pi_{b}(a|s)|da$
@@
    $\leq\frac{1}{\eta}\mathbb{E}_{s\sim d^{\pi_{\theta^{*}}}_{\rho}}\left(D_{\operatorname{KL}}(\pi_{\theta^{*}}(\cdot|s)\|\pi_{\theta_{t}}(\cdot|s))-D_{\operatorname{KL}}(\pi_{\theta^{*}}(\cdot|s)\|\pi_{\theta_{t+1}}(\cdot|s))\right)+\frac{2\eta M^{2}\left(1-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{r}^{t}\|^{2}}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}}$
@@
$\quad\sum_{t\in S_{\text{unsafe}}}\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)+\frac{1}{2}\sum_{t\in S_{\text{align}}}\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)+\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\cdot\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)$ (44)
@@
$e_{Q}$ $=\sum_{t\in S_{\text{safe}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\big{\|}Q_{r}^{\pi_{\theta_{t}}}-\hat{Q}_{r}^{\pi_{\theta_{t}}}\big{\|}_{2}+\sum_{t\in S_{\text{unsafe}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\big{\|}Q_{c}^{\pi_{\theta_{t}}}-\hat{Q}_{c}^{\pi_{\theta_{t}}}\big{\|}_{2}$
$+\sum_{t\in S_{\text{align}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\left[\frac{1}{2}\big{\|}Q_{r}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}^{\pi_{\theta_{t}}}_{r}(s,a)\big{\|}_{2}+\frac{1}{2}\big{\|}Q_{c}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{c}^{\pi_{\theta_{t}}}(s,a)\big{\|}_{2}\right]$ (45)
$+\sum_{t\in S_{\text{conflict}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\left[\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{r}^{t}\|^{2}}\right)\left\|Q_{r}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{r}^{\pi_{\theta_{t}}}(s,a)\right\|_{2}+\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\left\|Q_{c}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{c}^{\pi_{\theta_{t}}}(s,a)\right\|_{2}\right].$
Now, we need to upper bound the weighted critic error $e_{Q}$. We assume that there exists a positive constant $C$ such that $e_{Q}$ with $\frac{3CT(1+\eta M)}{(1-\gamma)^{2}}\epsilon_{\text{offline}}$ According to Gu et al. [2024b, Lemma A.6], by choosing reasonably large values for $h^{+}$ and $h^{-}$, we can ensure that:
@@
where the policy weights are assigned based on the categories of iterations:
  • Weight $w_{t}=$ $1$ for $t\in S_{\text{safe}}$,
  • Weight $w_{t}=$ 00 for $t\in S_{\text{unsafe}}$,
  • Weight $w_{t}=$ $\frac{1}{2}$ for $t\in S_{\text{align}}$,
  • Weight $w_{t}=$ $\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{\|g_{r}^{t}\|^{2}}$ for $t\in S_{\text{conflict}}$.
Under the weighted policy $\tilde{\pi}$, the following bound holds true for the reward value function:
@@
Now, by choosing $\eta=\sqrt{{\epsilon_{\text{dist}}(1-\gamma)^{3}}/{\big{(}{2\eta|\mathcal{S}||\mathcal{A}|M^{2}}+{3CT\eta M}(1-\gamma)\epsilon_{\text{offline}}}\big{)}}$, then we can ensure that
$V^{\pi^{*}}_{r}-V^{\tilde{\pi}}_{r}\leq\frac{1}{\sqrt{T}}\sqrt{\frac{32({2\eta|\mathcal{S}||\mathcal{A}|M^{2}}+{3(1+\eta M)}(1-\gamma)\epsilon_{\text{offline}})}{(1-\gamma)^{3}}}+\frac{3C\epsilon_{\text{offline}}}{(1-\gamma)^{2}}=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right)+\mathcal{O}(\epsilon_{\text{offline}}).$ (48)
@@
$V^{\tilde{\pi}}_{c}(\rho)-b$ $\leq\frac{4}{T}\left[\sum_{t\in S_{\text{safe}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)+\frac{1}{2}\sum_{t\in S_{\text{align}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)+\left(\frac{1}{2}-\frac{\langle g_{c}^{t},g_{r}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\sum_{t\in S_{\text{conflict}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)\right]$ (49)
@@
$\leq 2h^{+}+2\epsilon_{\text{offline}}=\mathcal{O}(\epsilon_{\text{offline}})+\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right).$

### v2

@@ 5.2 Ablation Study
Impact of Different Cost Limits. We evaluate our algorithm’s performance under varying cost limits $l=10,20$, and $30$, analyzing the learned policies’ behavior for each budget setting. The ablation results are presented in Figure 2 (b). Across all cost limit choices, our algorithm consistently achieves zero violations of the safety constraints, demonstrating its robustness to varying cost thresholds. This highlights the adaptability and reliability of our approach in maintaining safety compliance. While COptiDICE also strictly adheres to the cost limits, our method consistently outperforms it in terms of normalized return, demonstrating its superior ability to balance safety and reward optimization.
@@
Temperature Parameter Setting. We compare three diffusion-temperature schedules: (i) constant, $\beta_{t}=0.02$; (ii) square-root growth, $\beta_{t}=0.01\sqrt{t}$; and (iii) linear growth, $\beta_{t}=0.01\,t+0.04$, held fixed within each epoch. As shown in Figure 3 (Appendix C), all three schemes yield similar performance. Additional temperature ablations are reported in Appendix C.1.
Choice of Slack Variable. We introduce slack bounds so that reward maximization applies when $V_{\mathrm{norm}}\leq 1-h^{-}$ and cost minimization when $V_{\mathrm{norm}}\geq 1+h^{+}$. Both $h^{-}$ and $h^{+}$ are initialized to 0.2 and linearly decayed to zero; we ablate over initial values $h\in\{0.1,0.3,0.5,0.7\}$. Figure 4 (Appendix C) illustrates these results, with further slack-value studies in Appendix C.1.
Figure 2: (a) Computational Efficiency vs. Performance Trade-off. Normalized score ($y$-axis, combining reward and safety metrics) versus inference time for generating 1,000 actions ($x$-axis). (b) Normalized return and cost under varying cost limits ($l=10,20,30$). Since the cost is normalized relative to the corresponding cost limits, the safety threshold is consistently 1.0. The dashed line represents the safety boundary.
@@ 6 Related Work
Diffusion Models in Offline RL. Diffusion models’ powerful generative capacity for complex data has made them increasingly popular for modeling diverse offline datasets, valuable either for planning Janner et al. (2022); He et al. (2024); Ajay et al. (2022) or expressive policy modeling Chi et al. (2023); Wang et al. (2022); Fang et al. (2024); Hansen-Estruch et al. (2023); Chen et al. (2023); Lu et al. (2023). Their main drawback is inherently slow sampling due to many reverse-diffusion steps. To accelerate inference, methods like DDIM Song et al. (2020a) reduce steps via subsequence sampling, while DPM-Solver Lu et al. (2022) uses an optimized ODE solver to reach $\sim$10 steps. Useful extensions to safety-critical settings include the cost-constrained diffuser framework of Lin et al. (2023), SafeDiffuser Xiao et al. (2023) and OASIS’s conditional diffusion for dataset synthesis and careful distribution shaping Yao et al. (2024).
Online Safe RL. Safe RL in online environments has been extensively surveyed in Gu et al. (2024c). A common approach is constrained policy optimization—e.g., CPO Achiam et al. (2017) integrates TRPO Schulman (2015) for robust updates, and CRPO Xu et al. (2021) obviates complex dual-variable tuning. Primal–dual Lagrangian methods dynamically adjust multipliers on the fly to enforce safety criteria (Chow et al., 2018; Calian et al., 2020; Ding et al., 2020; Ying et al., 2022; Zhou et al., 2023), and more recent algorithms aim to better balance cost and reward via gradient manipulation Gu et al. (2024a) or refined trust-region formulations Kim et al. (2023, 2024). However, these methods depend on extensive environment interaction and accurate critic estimates, thus limiting their practicality under stringent real-world safety and data collection cost constraints.
Safe Offline RL. Considering safety specifically in the offline learning setting, Xu et al. (2022); Guan et al. (2023) penalized OOD and unsafe actions identified in the dataset. Diffusion-based safe policy learning was adapted by Lin et al. (2023) for cost-constrained scenarios, and decision-transformer architectures have been applied to safety considerations via Liu et al. (2023b). Energy-based diffusion policies can enforce hard constraints, for example, through weighted regression (Lu et al., 2023; Zheng et al., 2024; Koirala et al., 2024). More recent approaches learn by modelling conditional sequences as in Gong et al. (2025); Zhang et al. (2023) or by strategically sharing representations across constraints for improved adaptation Chemingui et al. (2025).
7 Conclusion and Limitations
We present DRCORL, a novel approach for offline constrained reinforcement learning that learns safe, high-performance policies from fixed datasets. DRCORL employs diffusion models to faithfully capture offline behavioral policies and distills them into simplified policies for fast inference. To balance reward maximization and constraint satisfaction, we utilize a gradient manipulation strategy that adapts dynamically without extensive hyperparameter tuning. Extensive experiments across standard safety benchmarks show that DRCORL consistently outperforms existing offline safe RL methods, achieving superior rewards while satisfying safety constraints.
We also identify several limitations offering avenues for future work. First, DRCORL’s pretraining stage comprising diffusion-model training and critic estimation, though producing lightweight policy at inference and avoids unsafe online exploration, incurs additional computational overhead. Future work should aim to optimize and parallelize the pretraining phase to further reduce its cost. Second, guaranteeing zero constraint violations remains challenging in offline RL, particularly when dataset quality and coverage vary. Developing methods with greater robustness to imperfect data and tighter safety guarantees is a promising direction. Despite these challenges, our contributions lay a strong foundation for advancing both generalization and safety in constrained offline reinforcement learning.
@@ Appendix A Proofs of Theoretical Results
$\left|\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\mathbb{E}_{a\sim\pi_{b}(\cdot|s)}[A^{\tilde{\pi}}_{c}(s,a)]\right|$ $\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\left|\int_{\mathcal{A}}\pi_{b}(a|s)Q^{\tilde{\pi}}(s,a)-\tilde{\pi}(a|s)Q^{\tilde{\pi}}(s,a)da\right|$
@@
$\leq\frac{c_{\max}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\tilde{\pi}}_{\rho}}\int_{\mathcal{A}}|\tilde{\pi}(a|s)-\pi_{b}(a|s)|da$ (28)
@@
We admit that ommitting the reverse KL divergence term simplifies the derivation of a closed-form NPG update, yet the core insights from this analysis are expected to remain relevant to the behavior of the full DRCORL algorithm. This is mainly because:
(i) The initial pretraining of the diffusion model (Section 3.1) and its subsequent use as a regularizer (Section 3.2) aim to ensure the learned policy $\pi_{\theta}$ primarily operates within a region close to the behavioral policy $\pi_{b}$. Assumption A.1 formalizes this by bounding the KL divergence.
(ii) In this regularized regime, the KL term primarily acts to constrain exploration and prevent significant deviation from the data distribution. The fundamental trade-offs and convergence dynamics related to balancing expected rewards and costs are still governed by the NPG updates on these value estimations. Thus, the analysis, while idealized, sheds light on the core learning dynamics concerning reward maximization under cost constraints.
@@
    $\leq\frac{1}{\eta}\mathbb{E}_{s\sim d^{\pi_{\theta^{*}}}_{\rho}}\left(D_{\operatorname{KL}}(\pi_{\theta^{*}}(\cdot|s)\|\pi_{\theta_{t}}(\cdot|s))-D_{\operatorname{KL}}(\pi_{\theta^{*}}(\cdot|s)\|\pi_{\theta_{t+1}}(\cdot|s))\right)$
    $\quad+\frac{2\eta M^{2}\left(1-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{r}^{t}\|^{2}}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}}$
@@
$\quad\sum_{t\in S_{\text{unsafe}}}\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)+\frac{1}{2}\sum_{t\in S_{\text{align}}}\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)$ (39)
$\quad\quad+\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\cdot\left(V_{c}^{\pi_{\theta^{*}}}(\rho)-V_{c}^{\pi_{\theta_{t}}}(\rho)\right)$
@@
$e_{Q}$ $=\sum_{t\in S_{\text{safe}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\big{\|}Q_{r}^{\pi_{\theta_{t}}}-\hat{Q}_{r}^{\pi_{\theta_{t}}}\big{\|}_{2}+\sum_{t\in S_{\text{unsafe}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\big{\|}Q_{c}^{\pi_{\theta_{t}}}-\hat{Q}_{c}^{\pi_{\theta_{t}}}\big{\|}_{2}$ (40)
$+\sum_{t\in S_{\text{align}}}\frac{3(1+\eta M)}{(1-\gamma)^{2}}\left[\frac{1}{2}\big{\|}Q_{r}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}^{\pi_{\theta_{t}}}_{r}(s,a)\big{\|}_{2}+\frac{1}{2}\big{\|}Q_{c}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{c}^{\pi_{\theta_{t}}}(s,a)\big{\|}_{2}\right]$
$+\sum_{t\in S_{\text{conflict}}}\hskip-7.0pt\frac{3(1+\eta M)}{(1-\gamma)^{2}}\bigg{[}\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{r}^{t}\|^{2}}\right)\hskip-3.0pt\left\|Q_{r}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{r}^{\pi_{\theta_{t}}}(s,a)\right\|_{2}$
$\hskip 85.35826pt+\left(\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\hskip-3.0pt\left\|Q_{c}^{\pi_{\theta_{t}}}(s,a)-\hat{Q}_{c}^{\pi_{\theta_{t}}}(s,a)\right\|_{2}\bigg{]}.$
Now, we need to upper bound the weighted critic error $e_{Q}$. We assume that there exists a positive constant $C$ such that $e_{Q}$ with $\frac{3CT(1+\eta M)}{(1-\gamma)^{2}}\epsilon_{\text{offline}}$ . According to Gu et al. [2024b, Lemma A.6], by choosing reasonably large values for $h^{+}$ and $h^{-}$, we can ensure that:
@@
where the policy weights are assigned based on the categories of iterations: 1. Weight $w_{t}=$ 11 for $t\in S_{\text{safe}}$, 2. Weight $w_{t}=$ 0 for $t\in S_{\text{unsafe}}$,
3. Weight $w_{t}=$ $\frac{1}{2}-\frac{\langle g_{r}^{t},g_{c}^{t}\rangle}{\|g_{r}^{t}\|^{2}}$ for $t\in S_{\text{conflict}}$. Under the weighted policy $\tilde{\pi}$, the following bound holds true for the reward value function:
@@
Now, by choosing $\eta=\sqrt{{\epsilon_{\text{dist}}(1-\gamma)^{3}}/{\big{(}{2\eta|\mathcal{S}||\mathcal{A}|M^{2}}+{3CT\eta M}(1-\gamma)\epsilon_{\text{offline}}}\big{)}}$, we can ensure that
$V^{\pi^{*}}_{r}-V^{\tilde{\pi}}_{r}$ $\leq\frac{1}{\sqrt{T}}\sqrt{\frac{32({2\eta|\mathcal{S}||\mathcal{A}|M^{2}}+{3(1+\eta M)}(1-\gamma)\epsilon_{\text{offline}})}{(1-\gamma)^{3}}}+\frac{3C\epsilon_{\text{offline}}}{(1-\gamma)^{2}}$ (43)
$=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right)+\mathcal{O}(\epsilon_{\text{offline}}).$
@@
$\quad\ V^{\tilde{\pi}}_{c}(\rho)-b$
$\leq\frac{4}{T}\left[\sum_{t\in S_{\text{safe}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)+\frac{1}{2}\sum_{t\in S_{\text{align}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)+\left(\frac{1}{2}-\frac{\langle g_{c}^{t},g_{r}^{t}\rangle}{2\|g_{c}^{t}\|^{2}}\right)\sum_{t\in S_{\text{conflict}}}(V^{\pi_{\theta_{t}}}_{c}(\rho)-b)\right]$
@@
$\leq 2h^{+}+2\epsilon_{\text{offline}}=\mathcal{O}(\epsilon_{\text{offline}})+\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right).$ (44)